# Burstiness and interpersonal foraging between human infants and caregivers in the vocal domain

### v1

@@ 1 Introduction
In this study, we propose a method to address this confound and more reliably test for the effect of adult vocal input on subsequent infant vocal behavior and vice versa, by controlling for the tendency for consecutive vocal IEIs to be positively correlated. We first calculate infant and adult caregiver vocal IEIs, distinguishing between IEIs based on whether the first event received a vocal response or not (see Fig. 1 and 27). We then assess the extent of temporal clustering in infants’ and adults’ spontaneous vocal productions, as evidenced by positive correlations between consecutive IEIs (Fig. 2A). Finally, we use IEI residuals from the correlation analysis to test whether adult responses predict shorter infant IEIs after controlling for infant vocal clustering (Fig. 2B) and whether infant responses predict shorter adult IEIs after controlling for adult vocal clustering. Vocal responses being predictive of shorter IEIs after controlling for correlations between consecutive IEIs would be consistent with the idea that adult vocal input following an infant vocalization serves as a reward, causing the infant to vocalize sooner than they otherwise would, and vice versa for the effect of infant vocal input on adult vocalizations. We used data from 10+ hour home audio recordings collected from infant-worn recorders when infants were 3, 6, 9, and 18 months old, enabling us to test for age effects for all measures mentioned above. Analyses were performed at the daylong and five-minute levels using automated sound segmentation and labeling and at the five-minute level using human annotations.
@@ 2.2 Social inputs are associated with shorter IEIs, controlling for previous IEI
Figure 4: Response effects on IEIs after controlling for previous IEI (A) Standardized regression coefficients ($\beta$; y-axis) from linear models testing the effect of receipt of an adult (Ad) response on the length of infant speech-related (ChSp) IEIs—after controlling for the positive correlations between previous and current IEIs (see Methods and Fig. 3A for details on the current IEI $\sim$ previous IEI control)—are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis), for LENA daylong data. 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p<0.001$) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Statistical analyses were performed separately for each infant age group: 3, 6, 9, and 18 months (indicated by line and circle color; see legend). IEIs associated with responses (Resp IEIs) were significantly shorter for all age groups for $T_{\rm resp}\geq 3$ s. (B) Standardized regression coefficients ($\beta$; y-axis scale shared with panel A) from linear models analogous to A are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis) for the three different 5-minute validation datasets (see legend): human listener-labeled 5 minute sections with all adult vocalizations included (H: All Adult; dark cyan), human listener-labeled 5 minute sections with only infant-directed adult vocalizations included (H: Child-directed Adult; green), and corresponding 5 minute sections as labeled by LENA (L: 5 min; maroon). 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p$ $<$ 0.001) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Validation data were not separated into age blocks. Unlike in panel A, $\beta$s for $T_{\rm{resp}}\geq 3$ s are not significant at our $p<0.001$ threshold except for the human-labeled validation dataset in which only infant-directed Ad vocalizations are included (see Supplementary Figs. S30 and S31 for significant $\beta$ values at less stringent significance thresholds). However, they do trend negative for all three validation datasets for $T_{\rm{resp}}\geq 3$ s. (C) $\beta$ values (y-axis) from a similar model as in A, testing the effect of infant response on adult IEI length are shown for LENA daylong data. Resp IEIs are significantly shorter for all age groups for $T_{\rm resp}\geq 2$ s. (D) $\beta$ values (y-axis shared with panel C) from a similar model as in B testing the effect of infant response on adult IEI length are shown for validation datasets. For human-labeled vocalizations, when only infant-directed adult vocalizations are included, there is a significant negative effect of infant response on adult IEI for 1 s $\geq T_{\rm{resp}}\leq 5$ s.
@@ 2.2.1 Effect of adult responses on infant IEIs
For $T_{\rm{resp}}\geq 2$ s, infant IEIs were significantly shorter ($p<0.001$) after receiving an adult response for LENA daylong data, as evidenced by the negative response $\beta$s in Fig. 4A. This pattern of negative $\beta$s holds for all age groups for full day LENA-segmented data, except for 6 month-old infants at $T_{\rm resp}=2$ s, in which case the trend was consistent but did not reach significance at $p<0.001$ (see Supplementary Fig. S30 for significant $\beta$ values at less stringent significance thresholds). For validation datasets using human-labeled 5-minute excerpts (Fig. 4B), the negative $\beta$ was significant only for the human-labeled dataset with only infant-directed adult vocalizations included at $T_{\rm resp}=3$ s. All other $\beta$s trended in a direction consistent with the negative $\beta$s found for the daylong LENA-segmented data for $T_{\rm{resp}}\geq 2$ s but were not statistically significant at the $p<0.001$ threshold (see Supplementary Fig. S31 for significant $\beta$ values at less stringent significance thresholds). The corresponding LENA-labeled 5-minute portions also had non-significant negative response $\beta$s for $T_{\rm{resp}}\geq 2$ s, suggesting that the non-significance may primarily be an issue of insufficient statistical power with these much smaller datasets.
@@
For $T_{\rm resp}<2$ s, IEIs were longer after an adult response for LENA daylong data, as evidenced by positive $\beta$ values for all age groups. This trend was significant at $p<0.001$ at $T_{\rm resp}=0.5$ s for infants at 6, 9, and 18 months, and at $T_{\rm resp}=1$ s for 6 month old infants. The tendency for adult vocalizations to be longer than most other sound types (see Supplementary Table S2 and Supplementary Fig. S2) may contribute to positive $\beta$ values at $T_{\rm resp}<2$ s, which correspond to cases where the shortest IEIs in the data are included in response analyses (see also Supplementary Fig. S27). IEIs associated with a response necessarily have an intervening adult vocalization. IEIs not associated with a response, on the other hand, can only have silence or sound types other than infant (speech and non-speech related) or adult—which may generally be shorter than adult vocalizations—intervening. At low response window thresholds, short IEIs for which the limiting factor on IEI length is the intervening sound type are more likely to be included in the analysis. As a result, for small $T_{\rm resp}$ values, Resp IEIs may tend to be longer than NoResp IEIs purely as a function of the duration of the intervening adult response vocalization. This pattern of positive $\beta$s for $T_{\rm resp}<2$ s also holds for the validation datasets, with significant $\beta$s at $T_{\rm resp}=0.5$ s for all three validation datasets and at $T_{\rm resp}=1$ s for the human-labeled 5-minute excerpts with only infant-directed adult vocalizations included. Thus, as long as the response window parameter, $T_{\rm{resp}}$, is sufficiently long, results are consistent with a tendency for adult vocal responses to promote infant vocalization, causing infants to vocalize again sooner than would be expected based on the infant’s previous inter-vocalization interval.
@@ 2.2.2 Effect of infant response on adult IEIs
Adult IEIs were shorter following an infant response for $T_{\rm{resp}}\geq 2$ s. This effect is significant for all infant age groups for the full-day LENA-labeled data (Fig. 4C). For the human-labeled 5-minute datasets (Fig. 4D), the response effect is significant for 1 s $\leq T_{\rm{resp}}\leq 5$ s when only infant-directed adult vocalizations are included in the analysis, with the significant effect sizes being stronger than the effect sizes for the human-labeled 5-minute datasets with all adult vocalizations included, indicating that infant responses affect infant-directed adult vocalizations more than non-infant-directed adult vocalizations. For human-labeled 5-minute portions with all adult vocalizations included and for LENA labels for those same 5-minute portions, response effects are significant for a more limited range of $T_{\rm{resp}}$ values.
@@ 3 Discussion
In conclusion, we believe this approach provides the strongest evidence to date for facilitatory effects of social engagement on both infant and adult vocal behavior, supporting the view that infant vocalization is an interpersonal foraging process. In the future, this approach could be used to compare child-caregiver vocal interactions across diverse cultural and socioeconomic contexts [69]. Moreover, the burstiness of communicative signals and the importance of social vocal engagement in infant vocal development are shared by a number of animal communication systems [70, 71, 72, 73, 74]. As such, the use of correlations between successive IEIs to control for temporal structure in behavioral event series as well as the interpersonal foraging framework proposed in this study could be applied to understand interactive behaviors in other animal species [75], such as marmosets, which also exhibit infant-caregiver vocal turn-taking [76], as well as in other communication modalities [16, 77], such as gestures in humans [78, 79, 80] and great apes [74].
@@ Acknowledgments
For detailed tables summarizing the effect of infant age on recording level response $\beta$s (none significant at $p<0.001$), see the OSF repository associated with this project.

### v2

@@ 1 Introduction
In this study, we propose a method to address this confound and more reliably test for the effect of adult vocal input on subsequent infant vocal behavior and vice versa, by controlling for the temporal clustering of infant and adult vocalizations. We first calculate infant and adult caregiver vocal IEIs, distinguishing between IEIs based on whether the first event received a vocal response or not (see Fig. 1 and 27). We then assess the extent of temporal clustering in infants’ and adults’ spontaneous vocal productions, as evidenced by positive correlations between consecutive IEIs (Fig. 2A). Finally, we use IEI residuals from the correlation analysis to test whether adult responses predict shorter infant IEIs after controlling for infant vocal clustering (Fig. 2B) and whether infant responses predict shorter adult IEIs after controlling for adult vocal clustering. Vocal responses being predictive of shorter IEIs after controlling for correlations between consecutive IEIs would be consistent with the idea that adult vocal input following an infant vocalization serves as a reward, causing the infant to vocalize sooner than they otherwise would, and vice versa for the effect of infant vocal input on adult vocalizations. We used data from 200 daylong (10+ hour) home audio recordings collected longitudinally from infant-worn recorders when infants were 3, 6, 9, and 18 months old (totaling over 2,400 hours of audio), enabling us to test for age effects for all measures mentioned above (see Methods and Supplementary Sections S2 and S3 for details of data collection and sample sizes). Analyses were performed at the daylong and five-minute levels using automated sound segmentation and labeling and at the five-minute level using human annotations.
@@ 2.2 Social inputs are associated with shorter IEIs, controlling for previous IEI
Figure 4: Response effects on IEIs after controlling for previous IEI (A) Standardized regression coefficients ($\beta$; y-axis) from linear models testing the effect of receipt of an adult (Ad) response on the length of infant speech-related (ChSp) IEIs—after controlling for the positive correlations between previous and current IEIs (see Methods and Fig. 3A for details on the current IEI $\sim$ previous IEI control)—are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis), for LENA daylong data. 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p<0.001$) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Statistical analyses were performed separately for each infant age group: 3, 6, 9, and 18 months (indicated by line and circle color; see legend). IEIs associated with responses (Resp IEIs) were significantly shorter for all age groups for $T_{\rm resp}\geq 3$ s. (B) Standardized regression coefficients ($\beta$; y-axis scale shared with panel A) from linear models analogous to A are shown as a function of the response window, $T_{\rm resp}$ (seconds; x-axis) for the three different 5-minute validation datasets (see legend): human listener-labeled 5 minute sections with all adult vocalizations included (H: All Adult; dark cyan), human listener-labeled 5 minute sections with only infant-directed adult vocalizations included (H: Child-directed Adult; green), and corresponding 5 minute sections as labeled by LENA (L: 5 min; maroon). 99.9% confidence intervals are shown as error bars and significant $\beta$ values (at $p$ $<$ 0.001) are indicated as solid circles. $\beta$ values are staggered around the relevant $T_{\rm resp}$ value for easier visualization. Validation data were not separated into age blocks. Unlike in panel A, $\beta$s for $T_{\rm{resp}}\geq 3$ s are not significant at our $p<0.001$ threshold except for the human-labeled validation dataset in which only infant-directed Ad vocalizations are included (see Supplementary Figs. S30 and S31 for significant $\beta$ values at less stringent significance thresholds). However, they do trend negative for all three validation datasets for $T_{\rm{resp}}\geq 3$ s. (C) $\beta$ values (y-axis) from a similar model as in A, testing the effect of infant response on adult IEI length are shown for LENA daylong data. Resp IEIs are significantly shorter for all age groups for $T_{\rm resp}\geq 2$ s. (D) $\beta$ values (y-axis shared with panel C) from a similar model as in B testing the effect of infant response on adult IEI length are shown for validation datasets. For human-labeled vocalizations, when only infant-directed adult vocalizations are included, there is a significant negative effect of infant response on adult IEI for 1 s $\leq T_{\rm{resp}}\leq 5$ s.
@@ 2.2.1 Effect of adult responses on infant IEIs
For $T_{\rm{resp}}\geq 2$ s, infant IEIs were significantly shorter ($p<0.001$) after receiving an adult response for LENA daylong data, as evidenced by the negative response $\beta$s in Fig. 4A. This pattern of negative $\beta$s holds for all age groups for full day LENA-segmented data, except for 6 month-old infants at $T_{\rm resp}=2$ s, in which case the trend was consistent but did not reach significance at $p<0.001$ (see Supplementary Fig. S30 for significant $\beta$ values at less stringent significance thresholds). For the representative case of $T_{\rm resp}=5$ s (see Fig. 2B), $\beta$s ranged from -0.09 to -0.15 across all age groups (see Supplementary Table S10 as well as the OSF repository associated with this study).
For validation datasets using human-labeled 5-minute excerpts (Fig. 4B), the negative $\beta$ was significant only for the human-labeled dataset with only infant-directed adult vocalizations included at $T_{\rm resp}=3$ s. All other $\beta$s trended in a direction consistent with the negative $\beta$s found for the daylong LENA-segmented data for $T_{\rm{resp}}\geq 2$ s but were not statistically significant at the $p<0.001$ threshold (see Supplementary Fig. S31 for significant $\beta$ values at less stringent significance thresholds and Supplementary Table S12 for $\beta$ values for the representative case of $T_{\rm resp}=5$ s). The corresponding LENA-labeled 5-minute portions also had non-significant negative response $\beta$s for $T_{\rm{resp}}\geq 2$ s, suggesting that the non-significance may primarily be an issue of insufficient statistical power with these much smaller datasets.
@@
For $T_{\rm resp}<2$ s, IEIs were longer after an adult response for LENA daylong data, as evidenced by positive $\beta$ values for all age groups. This trend was significant at $p<0.001$ at $T_{\rm resp}=0.5$ s for infants at 6, 9, and 18 months, and at $T_{\rm resp}=1$ s for 6 month old infants. The tendency for adult vocalizations to be longer than most other sound types (see Supplementary Table S2 and Supplementary Fig. S2) may contribute to positive $\beta$ values at $T_{\rm resp}<2$ s, which correspond to cases where the shortest IEIs in the data are included in response analyses (see also Supplementary Fig. S27). IEIs associated with a response necessarily have an intervening adult vocalization. IEIs not associated with a response, on the other hand, can only have silence or sound types other than infant (speech and non-speech related) or adult—which may generally be shorter than adult vocalizations—intervening. At low response window thresholds, short IEIs for which the limiting factor on IEI length is the intervening sound type are more likely to be included in the analysis. As a result, for small $T_{\rm resp}$ values, Resp IEIs may tend to be longer than NoResp IEIs purely as a function of the duration of the intervening adult response vocalization. This pattern of positive $\beta$s for $T_{\rm resp}<2$ s also holds for the validation datasets, with significant $\beta$s at $T_{\rm resp}=0.5$ s for all three validation datasets and at $T_{\rm resp}=1$ s for the human-labeled 5-minute excerpts with all adult vocalizations included. Thus, as long as the response window parameter, $T_{\rm{resp}}$, is sufficiently long, results are consistent with a tendency for adult vocal responses to promote infant vocalization, causing infants to vocalize again sooner than would be expected based on the infant’s previous inter-vocalization interval.
@@ 2.2.2 Effect of infant response on adult IEIs
Adult IEIs were shorter following an infant response for $T_{\rm{resp}}\geq 2$ s. This effect is significant for all infant age groups for the full-day LENA-labeled data (Fig. 4C). $\beta$s ranged from -0.11 to -0.17 across all infant age groups for the representative case of $T_{\rm resp}=5$ s (see Supplementary Table S11).
For the human-labeled 5-minute datasets (Fig. 4D), the response effect is significant for 1 s $\leq T_{\rm{resp}}\leq 5$ s when only infant-directed adult vocalizations are included in the analysis, with the significant effect sizes being stronger than the effect sizes for the human-labeled 5-minute datasets with all adult vocalizations included, indicating that infant responses affect infant-directed adult vocalizations more than non-infant-directed adult vocalizations. For human-labeled 5-minute portions with all adult vocalizations included and for LENA labels for those same 5-minute portions, response effects are significant for a more limited range of $T_{\rm{resp}}$ values.
@@ 3 Discussion
In the future, this approach could be used to compare child-caregiver vocal interactions across diverse cultural and socioeconomic contexts [69]. Moreover, the burstiness of communicative signals and the importance of social vocal engagement in infant vocal development are shared by a number of animal communication systems [70, 71, 72, 73, 74]. As such, the use of correlations between successive IEIs to control for temporal structure in behavioral event series as well as the interpersonal foraging framework proposed in this study could be applied to understand interactive behaviors in other animal species [75], such as marmosets, which also exhibit infant-caregiver vocal turn-taking [76], as well as in other communication modalities [16, 77], such as gestures in humans [78, 79, 80] and great apes [74].
In summary, the current study provides robust evidence for facilitatory effects of social engagement on both infant and adult vocalizations in naturalistic, ecologically relevant contexts, after controlling for the inherent burstiness of vocal behavior. These findings support the view that infant vocalization is an interpersonal foraging process.
@@ Acknowledgments
Tables with response effect $\beta$s using a response window ($T_{\rm resp}$) value of 5 s as a representative case are shown below. Response $\beta$s for the effect of adult (Ad) responses on infant speech-related (ChSp) IEIs and for the effect of infant speech-related responses on adult IEIs are shown for LENA daylong data and validation data. Results for the full range of $T_{\rm{resp}}$ values analyzed are available in the OSF repository associated with this study.
Infant age ChSp IEI Response $\beta$ 99.9% CI
3 months -0.12 (-0.16, -0.08)
6 months -0.09 (-0.13, -0.05)
9 months -0.11 (-0.15, -0.07)
18 months -0.15 (-0.19, -0.11)
Table S10: Effect of receiving an adult response on infant speech-related IEI length for LENA daylong data for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving an Ad response on ChSp IEI residuals (after regressing current IEI on previous IEI with infant ID as a random effect) for $T_{\rm resp}=5$ s are shown, by infant age. 99.9% confidence intervals for the $\beta$s are also provided. All effects are significant at the $p<0.001$ level.
Infant Age Ad IEI Response $\beta$ 99.9% CI
3 months -0.17 (-0.21, -0.12)
6 months -0.11 (-0.15, -0.06)
9 months -0.14 (-0.19, -0.10)
18 months -0.13 (-0.17, -0.09)
Table S11: Effect of receiving an infant response on adult IEI length for LENA daylong data for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving a ChSp response on Ad IEI residuals (after regressing current IEI on previous IEI with infant ID as a random effect) for $T_{\rm resp}=5$ s are shown, by infant age. 99.9% confidence intervals for the $\beta$s are also provided. All effects are significant at the $p<0.001$ level.
Data type ChSp IEI Response $\beta$ 99.9% CI
L (5min) -0.07 ($p$ = 0.05) (-0.18, 0.05)
H (5min; All Ad) -0.03 ($p$ = 0.30) (-0.14, 0.07)
H (5min; T-Ad) -0.07 ($p$ = 0.03) (-0.19, 0.04)
Table S12: Effect of receiving an adult response on infant speech-related IEI length for validation datasets for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving an Ad response on ChSp IEI residuals (after regressing current IEI on previous IEI with infant ID, infant age, and infant ID-infant age interaction as random effects) for $T_{\rm resp}=5$ s are shown for the validation datasets: human listener-labeled 5-minute sections with all adult vocalizations (H (5min; All-Ad)); human listener-labeled 5-minute sections with only child-directed adult vocalizations included (H (5min; T-Ad)); and the corresponding LENA subset of 5-minute sections (L (5min)). $p$ values are indicated in parentheses. Significant $\beta$s at $p$ = 0.05 are in bold (the $p$ value for L (5min) data is 0.0478 and has been rounded to 0.05). 99.9% confidence intervals for the $\beta$s are also provided.
Data type Ad IEI Response $\beta$ 99.9% CI
L (5min) -0.12 ($p$ = 0.002)** (-0.25, 0.01)
H (5min; All Ad) 0.04 ($p$ = 0.30) (-0.09, 0.17)
H (5min; T-Ad) -0.16*** ($p<$ 0.001) (-0.31, -0.001)
Table S13: Effect of receiving an infant response on adult IEI length for validation datasets for $T_{\rm resp}=5$ s. Standardized regression coefficients ($\beta$s) from linear models testing the effect of receiving a ChSp response on Ad IEI residuals (after regressing current IEI on previous IEI with infant ID, infant age, and infant ID-infant age interaction as random effects) for $T_{\rm resp}=5$ s are shown for the validation datasets. $p$ values are indicated in parentheses. Significant $\beta$s at $p$ = 0.05 are in bold, significant $\beta$s at $p$ = 0.01 are indicated by **, and significant $\beta$s at $p$ = 0.001 are indicated by ***. 99.9% confidence intervals for the $\beta$s are also provided.
For tables summarizing the effect of infant age on recording level response $\beta$s (none significant at $p<0.001$), see the OSF repository associated with this project.